OpenStudy (anonymous):

A torque of 60 N*m acts on a wheel of moment of inertia 30 kg*m^2 for 5 s and then is removed. a) What is the angular acceleration of the wheel? b) How many revolutions does it make in 15 s if it starts at rest?

OpenStudy (anonymous):

\[\tau=I a\]

OpenStudy (anonymous):

\[60=30 * a\]

OpenStudy (anonymous):

k thx and wat about b) which formula do i apply?

OpenStudy (anonymous):

You have the angular acceleration which is:\[\alpha =d^2\theta/dt^2\]so, to figure out how many revolutions have happened in 15s we integrate twice to get the total angle swept out:\[d^2\theta=\alpha dt\]and,\[\Delta \theta=\int\limits_{0}^{15}2dt=30\]One revolution represents\[\Delta \theta=2\pi\]so 30 radians represents:\[\frac{30}{2\pi}\approx4.77\]So about 4.77 revolutions have occurred in 15s, starting from rest.

OpenStudy (anonymous):

ah crap I only integrated once here. We need to integrate twice! lol

OpenStudy (anonymous):

\[\Delta \theta=15^2=225\]So we get\[\frac{225}{2\pi}\approx35.8\]so 35.8 revolutions should be correct :)

OpenStudy (anonymous):

i tried 35.8 but apparently its wrong

OpenStudy (anonymous):

5sec , not 15 sec

OpenStudy (anonymous):

ooops...the force only acts for 5 seconds not 15! so for 10 seconds the angular acceleration is zero (for the first 5s the angular acceleration is 2)

OpenStudy (anonymous):

gotta go tho...this should be enough for you to solve it :)

OpenStudy (anonymous):

ok, in case you still needed help on this one:\[\theta=\frac{1}{2}\alpha t^2+\omega t+\theta _{0}\]This is the general equation of motion you need. Setting initial angular position=0 and angular acceleration=2:\[\theta=5^2+2(10)+0=45\]So total number of revolutions is:\[\frac{45}{2\pi}\approx 7.2\]

OpenStudy (anonymous):

i got 19.894

OpenStudy (anonymous):

yeah I think I see my error

OpenStudy (anonymous):

5^2+2(5)(10)+0=125 and 125/2pi=19.9

OpenStudy (anonymous):

good call

OpenStudy (anonymous):

thank you though

OpenStudy (anonymous):

np...im a little rusty hehe